# MLLM-CL: Continual Learning for Multimodal Large Language Models

### v1

@@ A.1 Training Details
Sec. A.1 shows the hyperparameters for training the router and expert in domain continual learning. For most configurations, we follow the default setting of LLaVA 1.5 liu2023llava . In terms of the training epoch, we ensure that each task receives comparable training exposure across datasets of varying sizes, equivalent to around 60,000 training instances. For all the experiments, we use 8 A100 GPUs and the training time for each task is around 1.5 hours. Sec. A.1 shows the hyperparameters for ability continual learning. For ability continual learning, the training time is around 20 hours to train all the tasks sequentially. For the router training, we use a small learning rate of 5e-6 and train 40 epochs in domain continual learning and 35 epochs for ability continual learning. The batch size is set to 32, with other configurations identical to those of each expert. We use the codebase from CoIN chen2024coin and LLaVA liu2023llava .
Table 7: Hyperparameters of MR-LoRA in domain continual learning.
Config Value
optimizer AdamW
batch size 64
lr schedule cosine decay
lr warmup ratio 0.03
base lr 2e-5
epoch for RS 1
epoch for Med 3
epoch for AD 1
epoch for Sci 2
epoch for Fin 1
LoRA rank 128
DeepSpeed stage 2
Table 8: Hyperparameters of MR-LoRA in ability continual learning.
Config Value
optimizer AdamW
batch size 64
lr schedule cosine decay
lr warmup ratio 0.03
base lr (M&L, VP, GUI) 2e-5
base lr (OCR) 5e-6
epoch for OCR 3
epoch for Math & Logic 1
epoch for VP 1
epoch for GUI agent 3
LoRA rank 128
DeepSpeed stage 2
@@ A.2 Evaluation Details
In domain continual learning, for the financial task, all the questions are MCQ or Y/N questions; we require the prediction to exactly match the ground truth. For autonomous driving, medical and remote sensing tasks, we regard the prediction should includes the ground truth as the correct answer. This serves as the default evaluation method. For science tasks, some test samples are multiple-choice questions (MCQs), and predictions are required to exactly match the ground truth. Certain questions in MapQA chang2022mapqa require the model to list places; in these cases, we compute the percentage of correct responses. Other science questions are evaluated according to the default method. In ability continual learning, we follow the default setting of the corresponding benchmarks.
@@ Appendix B Details of StockQA Dataset
Overview. StockQA dataset is a multimodal financial dataset concentrated on stock analysis. It is rewritten from FinVis wang2023finvis dataset. Finvis dataset is a Chinese caption dataset generated by GPT4V achiam2023gpt . All the captions are related to the stock technical indicator analysis. However, caption form is not convenient for evaluation and there may be a language gap between this task and other tasks. Therefore, we use a questioner-inspector data pipeline using a powerful MLLM Qwen2.5-VL Qwen2.5-VL to rewrite the caption into MCQ and Y/N question-answer pairs and name it StockQA. During the checking process for the inspector, we find that the inspector misclassified some correct question-answer pairs, but it can find some wrong cases and ensure the correctness of the final dataset.
Prompts for agents. Figs. 8 and 9 shows the prompt we use for the Questioner to generate Y/N and MCQ question-answer pairs, respectively. Fig. 10 is the prompt we use for the inspector.
Rules for filtering. After using an inspector agent to check the correctness and rationality, we use the following rules to balance the choices of multiple choice questions to avoid the position bias liu2024lost and format the output.
@@
Statistics of StockQA dataset. StockQA is a new VQA dataset related to multimodal stock analysis. It includes 70k question-answer pairs. of which 60k is the training set and 10k is the test set. For the training data, there are 40k MCQ and 20k Y/N QA pairs. For the test data, there are 8k MCQ and 2k QA pairs. Each choice is equally distributed after our cleaning process. Figs. 13 and 12 shows the word cloud and examples of StockQA dataset. Tabs. 9 and 11 shows the detailed statistics of StockQA dataset.
Dataset License. Our dataset follows the CC-BY license. This license allows reusers to distribute, remix, adapt, and build upon the material in any medium or format, so long as attribution is given to the creator. The license allows for commercial use. For other assets we used, we list the licenses below in Tab. 10.
Figure 10: Prompt for the Inspector to check the question answer pairs.
Table 9: Statistics of StockQA dataset.
@@
CC-BY-4.0 CLEVR, ScreenQA, Screen2Words, MMTBench
@@
Figure 14: Prompt for the router in ability continual learning.
@@ Appendix C Detailed Continual Learning Results
In this section, we show the detailed inference results of all the methods (LoRA hu2021lora , LoRA hu2021lora , MoELoRA chen2024coin , MoELoRA chen2024coin , L2P l2p , L2P l2p , ModalPrompt zeng2024modalprompt , ModalPrompt zeng2024modalprompt , HiDe-LLaVA guo2025hide , HiDe-LLaVA guo2025hide and MR-LoRA) during each continual learning stage, where  denotes the original method with replay data.
@@ C.1 Baseline Results in Domain Continual Learning
Table 11: Result matrix of LoRA.
RS Med AD Sci Fin
RS 80.72
Med 56.10 63.25
AD 34.96 34.75 51.81
Sci 26.20 19.60 3.13 25.62
Fin 26.75 25.76 0.79 18.69 70.44
Table 12: Result matrix of LoRA.
RS Med AD Sci Fin
RS 80.87
Med 55.85 61.35
AD 34.53 30.27 52.04
Sci 56.00 36.95 27.09 25.61
Fin 42.27 29.32 22.91 21.67 67.71
Table 13: Result matrix of MoELoRA.
RS Med AD Sci Fin
RS 80.05
Med 55.74 58.78
AD 29.98 29.43 51.69
Sci 38.48 13.70 0.83 25.58
Fin 21.42 25.29 0.79 17.01 60.34
Table 14: Result matrix of MoELoRA.
RS Med AD Sci Fin
RS 80.00
Med 53.61 60.20
AD 27.50 24.64 51.92
Sci 39.42 30.61 30.96 25.79
Fin 47.91 24.27 17.63 17.27 49.71
Table 15: Result matrix of O-LoRA.
RS Med AD Sci Fin
RS 80.22
Med 77.55 57.57
AD 66.30 37.21 51.02
Sci 64.38 37.71 28.40 46.64
Fin 62.68 35.17 16.93 34.44 92.16
Table 16: Result matrix of O-LoRA.
RS Med AD Sci Fin
RS 80.13
Med 77.05 63.40
AD 73.45 59.08 51.53
Sci 75.27 54.81 36.47 46.79
Fin 76.21 51.34 36.50 42.64 90.20
Table 17: Result matrix of L2P.
RS Med AD Sci Fin
RS 80.02
Med 77.42 60.30
AD 68.99 33.84 51.89
Sci 64.55 38.90 27.40 49.61
Fin 63.82 34.63 22.96 38.58 92.98
Table 18: Result matrix of L2P.
RS Med AD Sci Fin
RS 80.09
Med 78.04 59.23
AD 73.05 38.72 52.60
Sci 75.52 42.19 27.40 49.61
Fin 75.21 38.50 32.31 41.05 88.05
Table 19: Result matrix of ModalPrompt.
RS Med AD Sci Fin
RS 80.11
Med 69.22 50.10
AD 66.41 40.33 33.84
Sci 68.60 38.77 24.26 40.23
Fin 65.99 37.35 23.27 37.61 87.60
Table 20: Result matrix of ModalPrompt.
RS Med AD Sci Fin
RS 80.11
Med 68.87 53.10
AD 64.51 42.08 45.42
Sci 64.50 39.37 26.21 37.81
Fin 64.77 38.60 20.61 29.98 88.22
Table 21: Result matrix of HiDe-LLaVA.
RS Med AD Sci Fin
RS 80.91
Med 70.48 60.46
AD 33.70 33.78 51.87
Sci 38.78 33.15 14.76 44.97
Fin 41.17 30.33 18.73 37.08 92.21
Table 22: Result matrix of HiDe-LLaVA.
RS Med AD Sci Fin
RS 81.51
Med 76.89 47.84
AD 72.32 42.51 33.28
Sci 74.99 42.01 37.89 47.53
Fin 75.36 39.23 37.17 45.02 81.89
@@ C.2 Baseline Results in Ability Continual Learning
Table 23: Result matrix of LoRA.
OCR Math VP GA
OCR 31.10
Math 2.10 28.40
VP 29.00 29.00 60.15
GA 10.80 26.40 27.16 30.83
Table 24: Result matrix of LoRA.
OCR Math VP GA
OCR 31.10
Math 3.10 31.30
VP 29.10 29.30 61.68
GA 12.00 22.50 14.34 33.33
Table 25: Result matrix of MoELoRA.
OCR Math VP GA
OCR 30.60
Math 2.10 30.70
VP 29.50 27.40 59.39
GA 15.10 28.30 11.80 27.78
Table 26: Result matrix of MoELoRA.
OCR Math VP GA
OCR 30.60
Math 4.30 27.30
VP 28.60 29.10 58.63
GA 11.90 24.70 23.10 23.75
Table 27: Result matrix of O-LoRA.
OCR Math VP GA
@@
Math 33.20 32.50
VP 33.00 33.50 65.99
GA 18.50 26.10 51.90 32.50
Table 28: Result matrix of O-LoRA.
OCR Math VP GA
@@
Math 33.40 32.50
VP 33.00 32.40 65.61
GA 28.60 32.80 51.40 32.50
Table 29: Result matrix of L2P.
OCR Math VP GA
OCR 31.40
Math 31.30 32.20
VP 30.40 32.30 62.06
GA 25.10 32.00 48.22 16.25
Table 30: Result matrix of L2P.
OCR Math VP GA
OCR 32.30
Math 28.30 31.20
VP 29.50 25.60 65.23
GA 25.50 26.10 61.80 18.47
Table 31: Result matrix of ModalPrompt.
OCR Math VP GA
OCR 31.00
Math 31.00 25.40
VP 31.00 32.80 61.68
GA 31.00 31.80 61.68 23.75
Table 32: Result matrix of ModalPrompt.
OCR Math VP GA
OCR 31.00
Math 31.00 24.40
VP 31.00 24.40 60.91
GA 31.00 24.40 60.91 21.25
Table 33: Result matrix of HiDe-LLaVA.
OCR Math VP GA
@@
Math 27.50 37.40
VP 21.90 29.00 61.29
GA 21.70 30.10 63.96 25.00
Table 34: Result matrix of HiDe-LLaVA.
OCR Math VP GA
OCR 30.40
Math 27.70 40.10
VP 26.60 33.80 61.17
GA 24.20 27.20 44.16 34.17

### v2

@@ A.1 Training Details
DCL. Tab.˜9 shows the hyperparameters for training the router and expert in domain continual learning. For most configurations, we follow the default setting of LLaVA 1.5 (Liu et al., 2023). To ensure comparable training exposure across datasets of varying sizes, each task is trained for approximately 60,000 instances in DCL. For efficient fine-tuning, a rank of 32 is employed. For all the experiments, we use 8 A100 GPUs, and the training time for each task is around 1 hour.
ACL. Tab.˜10 shows the hyperparameters for ability continual learning. For ability continual learning, training time is around 20 hours to train all the tasks sequentially.
Router Training. For the router training, we train 30 epochs in domain continual learning and ability continual learning; we keep other configurations identical to the experts’ except for the learning rate. We use the codebase from MCITlib (Guo et al., 2025d) and LLaVA (Liu et al., 2023).
Table 9: Hyperparameters of MR-LoRA in domain continual learning
Expert Config Router Config
LLaVA InternVL LLaVA InternVL
optimizer AdamW AdamW
batch size $64$ $64$
lr schedule cosine decay cosine decay
lr warmup ratio $0.03$ $0.03$
LoRA rank $32$ $32$
DeepSpeed stage 22 22
base lr $1\text{\times}{10}^{-4}$ $2\text{\times}{10}^{-5}$ $1\text{\times}{10}^{-4}$
epoch for RS 11 -
epoch for Med 33 $30$
epoch for AD 11 $30$
epoch for Sci 22 $30$
epoch for Fin 11 $30$
@@ A.2 Evaluation Details
In domain continual learning, for the financial task, all the questions are MCQ or Y/N questions; we require the prediction to exactly match the ground truth. For autonomous driving, medical, and remote sensing tasks, we consider the prediction to include the ground truth as the correct answer. This serves as the default evaluation method. For science tasks, some test samples are multiple-choice questions (MCQs), and predictions are required to exactly match the ground truth. Certain questions in MapQA (Chang et al., 2022) require the model to list places; in these cases, we compute the percentage of correct responses. Other science questions are evaluated according to the default method. In ability continual learning, we follow the default setting of the corresponding benchmarks.
A.3 Detailed Evalution Metrics
We used the integrated metrics in SEFE and MCITlib (Chen et al., 2025; Guo et al., 2025d) to evaluate the performance of each method.
  • Last accuracy is the accuracy of all seen tasks after learning the last task.
  • Mean Finetune Accuracy (MFT) measures the average accuracy achieved on each task immediately after it is learned, serving as an upper bound that reflects the model’s performance in the absence of forgetting.
  • Mean Final Accuracy (MFN) computes the average accuracy over all tasks after completing the full incremental training process, representing the model’s overall retained performance.
  • Mean Average Accuracy (MAA) calculates the mean of average accuracies on all learned tasks after each training step, offering a holistic view of performance throughout the continual learning process.
  • Backward Transfer (BWT) captures the change in accuracy for each task by comparing its final accuracy with that immediately after it was learned, quantifying the extent of forgetting.
For clarity, a conceptual illustration of the evaluation metrics is provided in  Fig.˜7.
Table 10: Hyperparameters of MR-LoRA in ability continual learning
Expert Config Router Config
LLaVA InternVL LLaVA InternVL
optimizer AdamW AdamW
batch size $128$ $128$
lr schedule cosine decay cosine decay
lr warmup ratio $0.03$ $0.03$
LoRA rank $32$ $32$
DeepSpeed stage 22 22
base lr (OCR) $5\text{\times}{10}^{-5}$ $2\text{\times}{10}^{-4}$ - -
base lr (M&L,VP,GUI) $2\text{\times}{10}^{-4}$ $2\text{\times}{10}^{-4}$ $2\text{\times}{10}^{-4}$ $1\text{\times}{10}^{-4}$
epoch for OCR 33 -
epoch for Math & Logic 11 $30$
epoch for VP 11 $30$
epoch for GUI Agent 33 $30$
Figure 7: Illustration of the evaluation metric calculations (Guo et al., 2025d).
A.4 Router prompt for MR-LoRA
We previously provided our router prompt for DCL in Fig.˜3. The prompt for ACL appears in Fig.˜8.
Figure 8: Prompt for the router in ability continual learning.
Figure 11: Prompt for the Inspector to check the question answer pairs.
@@ Appendix B Details of StockQA Dataset
Overview. The StockQA dataset is a multimodal financial dataset concentrated on stock analysis. It is rewritten from the FinVis (Wang et al., 2023b) dataset.
Finvis dataset is a Chinese caption dataset generated by GPT4V (Achiam et al., 2023). All the captions are related to the stock technical indicator analysis. However, the caption form is not convenient for evaluation, and there may be a language gap between this task and other tasks. Therefore, we use a questioner-inspector data pipeline with a powerful MLLM Qwen2.5-VL (Bai et al., 2025) to rewrite the caption into MCQ and Y/N question-answer pairs and name it StockQA. When manually checking the inspector process, we find that the inspector misclassified some correct question-answer pairs. Nevertheless, it successfully identified erroneous instances, thereby contributing to the overall correctness of the final dataset.
Prompts for agents. Figs.˜10 and 9 shows the prompt we use for the Questioner to generate Y/N and MCQ question-answer pairs, respectively. Fig.˜11 is the prompt we use for the inspector.
Rules for filtering. After using an inspector agent to check the correctness and rationality, we employ the following rules to balance the choices of multiple choice questions to mitigate the position bias (Liu et al., 2024c) and format the output.
@@
Table 11: Statistics of the StockQA dataset.
@@
$C$-BY-4.0 CLEVR, ScreenQA, Screen2Words, MMTBench
@@
Statistics of StockQA dataset. StockQA is a new VQA dataset related to multimodal stock analysis. It includes 70k question-answer pairs. of which 60k is the training set and 10k is the test set. For the training data, there are 40k MCQ and 20k Y/N QA pairs. For the test data, there are 8k MCQ and 2k QA pairs. Each choice is equally distributed after our cleaning process. Figs.˜14 and 13 shows the word cloud and examples of StockQA dataset. Tabs.˜11 and 12 shows the detailed statistics of StockQA dataset.
Dataset License. Our dataset follows the CC-BY license. This license allows reusers to distribute, remix, adapt, and build upon the material in any medium or format, so long as attribution is given to the creator. The license allows for commercial use. For other assets we used, we list the licenses below in Tab.˜12.
@@ Appendix C Detailed Continual Learning Results
In this section, we show the detailed inference results of all the methods (LoRA (Hu et al., 2021), LoRA (Hu et al., 2021), O-LoRA (Wang et al., 2023a), O-LoRA (Wang et al., 2023a), MoELoRA (Chen et al., 2024a), MoELoRA (Chen et al., 2024a), CL-MoE (Huai et al., 2025), CL-MoE (Huai et al., 2025), HiDe (Guo et al., 2025a), HiDe (Guo et al., 2025a), SEFE (Chen et al., 2025), SEFE (Chen et al., 2025), DISCO (Guo et al., 2025b), DISCO (Guo et al., 2025b) and MR-LoRA) during each continual learning stage, where  denotes the original method with replay data.
@@ C.1 Baseline Results in Domain Continual Learning
Table 13: Result matrices of InternVL-based baselines in domain continual learning. denotes the original method with replay data.
LoRA-FT RS Med AD Sci Fin
RS 81.29
Med 75.71 65.92
AD 69.38 56.87 53.56
Sci 71.12 53.75 46.83 53.48
Fin 69.93 52.17 33.04 42.67 91.07
LoRA-FT RS Med AD Sci Fin
RS 81.68
Med 77.45 66.69
AD 77.24 61.32 53.81
Sci 77.89 55.43 49.13 53.53
Fin 77.06 47.55 42.67 43.31 91.44
MoELoRA RS Med AD Sci Fin
RS 81.22
Med 77.56 66.00
AD 74.56 58.74 53.62
Sci 72.62 54.77 47.65 52.75
Fin 69.90 52.08 33.17 42.19 90.58
MoELoRA RS Med AD Sci Fin
RS 80.75
Med 78.10 64.77
AD 73.24 59.54 52.90
Sci 76.82 53.64 42.11 51.24
Fin 76.74 52.65 38.81 42.15 89.84
HiDe RS Med AD Sci Fin
RS 81.24
Med 79.59 64.71
AD 78.85 58.37 41.75
Sci 78.33 58.51 39.94 49.99
Fin 75.40 57.66 36.73 41.48 88.59
HiDe RS Med AD Sci Fin
RS 73.92
Med 71.44 64.22
AD 65.74 49.83 42.92
Sci 70.64 54.79 40.47 50.75
Fin 53.17 52.61 40.85 47.04 89.17
DISCO RS Med AD Sci Fin
RS 81.42
Med 79.13 63.80
AD 78.62 60.79 53.98
Sci 77.40 52.21 53.74 54.18
Fin 75.12 50.69 52.41 50.67 90.86
DISCO RS Med AD Sci Fin
RS 81.49
Med 80.14 63.05
AD 78.87 57.42 53.77
Sci 78.67 52.80 53.56 53.52
Fin 77.90 47.50 49.13 49.37 90.92
Table 14: Result matrices of LLaVA-based baselines in domain continual learning. denotes the original method with replay data.
LoRA-FT RS Med AD Sci Fin
RS 78.32
Med 74.68 57.53
AD 68.93 47.19 52.15
Sci 75.12 45.56 38.46 49.44
Fin 69.65 41.59 25.43 40.88 87.45
LoRA-FT RS Med AD Sci Fin
RS 79.33
Med 76.45 57.58
AD 74.54 54.26 52.96
Sci 77.00 50.31 45.13 51.88
Fin 76.54 50.27 43.01 43.32 89.85
O-LoRA RS Med AD Sci Fin
RS 79.25
Med 74.05 56.52
AD 76.06 43.71 52.32
Sci 76.60 44.87 40.57 50.58
Fin 74.64 44.42 30.02 41.47 87.15
O-LoRA RS Med AD Sci Fin
RS 79.17
Med 78.21 56.65
AD 77.52 38.60 37.81
Sci 77.61 44.22 35.40 45.59
Fin 76.94 41.17 34.18 39.61 83.22
MoELoRA RS Med AD Sci Fin
RS 79.09
Med 74.78 58.73
AD 77.69 43.72 51.47
Sci 76.87 43.79 32.81 48.67
Fin 77.54 41.85 27.62 40.13 86.75
MoELoRA RS Med AD Sci Fin
RS 79.66
Med 78.44 60.50
AD 78.54 49.86 52.54
Sci 78.00 50.53 43.32 49.30
Fin 77.63 49.54 39.08 41.04 89.21
CL-MoE RS Med AD Sci Fin
RS 79.08
Med 73.48 60.56
AD 72.61 44.42 51.62
Sci 71.02 48.04 37.70 50.28
Fin 71.34 46.84 26.33 41.17 88.74
CL-MoE RS Med AD Sci Fin
RS 79.40
Med 76.32 61.10
AD 72.01 54.49 52.56
Sci 76.64 53.89 43.83 49.98
Fin 76.58 52.31 39.65 45.64 90.21
HiDe RS Med AD Sci Fin
RS 78.14
Med 74.26 58.05
AD 74.90 42.94 39.65
Sci 75.43 44.91 38.33 46.44
Fin 74.31 48.95 33.21 38.54 81.55
HiDe RS Med AD Sci Fin
RS 79.21
Med 77.79 60.88
AD 77.64 48.19 38.12
Sci 77.51 48.84 35.76 46.71
Fin 74.80 42.29 34.03 38.01 79.22
SEFE RS Med AD Sci Fin
RS 78.27
Med 76.32 58.42
AD 77.22 49.13 52.49
Sci 77.83 47.70 43.01 49.04
Fin 77.26 50.37 37.21 40.87 86.82
SEFE RS Med AD Sci Fin
RS 79.21
Med 78.39 60.93
AD 79.00 57.68 53.11
Sci 78.76 51.39 47.99 51.87
Fin 78.43 52.85 46.21 47.76 89.33
DISCO RS Med AD Sci Fin
RS 78.57
Med 75.80 52.36
AD 76.37 49.78 53.04
Sci 76.11 45.76 49.26 49.23
Fin 76.03 45.20 43.79 42.33 88.95
DISCO RS Med AD Sci Fin
RS 79.20
Med 77.96 52.44
AD 78.05 49.85 53.03
Sci 77.26 46.32 53.08 51.99
Fin 77.78 46.25 50.45 49.51 89.71
@@ C.2 Baseline Results in Ability Continual Learning
Table 15: Result matrices of LLaVA-based baselines in ability continual learning. denotes the original method with replay data.
LoRA-FT OCR Math VP APP
OCR 33.30
Math 32.60 34.20
VP 31.70 32.80 65.10
APP 23.60 33.70 55.84 32.50
LoRA-FT OCR Math VP APP
OCR 32.60
Math 33.60 33.80
VP 31.10 33.50 66.12
APP 21.80 32.70 58.38 28.75
O-LoRA OCR Math VP APP
@@
Math 29.80 33.60
VP 27.40 33.70 58.63
APP 29.60 32.90 52.41 33.75
O-LoRA OCR Math VP APP
OCR 34.00
Math 28.40 36.80
VP 28.90 33.90 61.55
APP 29.60 31.30 60.79 27.50
MoELoRA OCR Math VP APP
OCR 32.70
Math 32.50 33.30
VP 30.80 33.00 64.59
APP 26.70 32.80 56.85 27.22
MoELoRA OCR Math VP APP
OCR 32.70
Math 29.40 33.10
VP 32.60 32.50 65.61
APP 19.80 32.20 54.19 30.00
CL-MoE OCR Math VP APP
OCR 33.00
Math 32.30 33.60
VP 30.20 32.50 64.72
APP 19.90 32.70 53.43 30.69
CL-MoE OCR Math VP APP
OCR 33.20
Math 34.30 36.70
VP 32.00 33.20 64.97
APP 25.40 31.80 60.91 30.00
HiDe OCR Math VP APP
OCR 33.40
Math 30.90 32.80
VP 30.40 33.30 56.98
APP 24.60 32.10 46.32 28.75
HiDe OCR Math VP APP
OCR 34.10
Math 32.60 35.70
VP 30.70 32.60 53.81
APP 24.60 28.40 30.71 23.75
SEFE OCR Math VP APP
OCR 33.00
Math 32.20 32.60
VP 31.80 33.30 64.59
APP 26.00 33.40 57.74 33.75
SEFE OCR Math VP APP
OCR 33.60
Math 33.80 37.50
VP 32.80 36.10 66.50
APP 25.60 34.80 57.61 31.39
DISCO OCR Math VP APP
@@
Math 31.80 33.40
VP 31.00 34.50 59.64
APP 32.90 33.10 60.15 30.14
DISCO OCR Math VP APP
OCR 33.40
Math 32.10 36.60
VP 32.20 37.00 63.07
APP 34.20 35.00 61.55 27.50
Table 16: Result matrices of InternVL-based baselines in ability continual learning. denotes the original method with replay data.
LoRA-FT OCR Math VP APP
OCR 32.20
Math 33.10 33.30
VP 31.80 32.30 68.02
APP 21.40 32.80 60.28 29.86
LoRA-FT OCR Math VP APP
@@
Math 35.30 35.40
VP 32.60 31.10 68.27
APP 26.30 34.20 62.56 31.25
O-LoRA OCR Math VP APP
OCR 32.70
Math 31.10 34.20
VP 30.20 33.00 63.20
APP 25.50 32.30 64.59 24.44
O-LoRA OCR Math VP APP
OCR 34.00
Math 30.90 34.40
VP 31.00 33.20 65.86
APP 21.70 31.10 59.77 31.25
MoELoRA OCR Math VP APP
OCR 32.20
Math 29.90 33.30
VP 29.20 32.80 67.64
APP 17.20 32.70 55.33 32.50
MoELoRA OCR Math VP APP
OCR 32.90
Math 31.50 36.50
VP 30.90 32.30 65.74
APP 13.90 29.70 54.95 32.50
HiDe OCR Math VP APP
OCR 33.40
Math 26.30 33.60
VP 30.10 33.00 61.80
APP 17.70 33.00 41.12 20.28
HiDe OCR Math VP APP
OCR 33.40
Math 28.10 34.70
VP 31.10 32.20 55.33
APP 25.30 29.20 42.13 20.28
DISCO OCR Math VP APP
OCR 31.90
Math 31.70 34.00
VP 32.10 33.50 63.45
APP 30.60 33.10 65.36 27.50
DISCO OCR Math VP APP
OCR 34.70
Math 31.50 34.70
VP 31.50 34.60 62.31
APP 32.30 32.30 64.97 30.14